OpenStudy (anonymous):

The system of equations is coincident. What are the missing values? 3x + y = 7 9x +___ y=____

OpenStudy (anonymous):

@0487308

OpenStudy (0487308):

Did the question tell you the solutions to x and y?

OpenStudy (anonymous):

nope

OpenStudy (lightn1ng):

How did they get 9x from 7?

OpenStudy (anonymous):

idk bruh

OpenStudy (0487308):

9x + 3y = 21

OpenStudy (0487308):

That is the solution.

OpenStudy (anonymous):

so my first answer is 21

OpenStudy (0487308):

The blanks in order are 3 and 21.

OpenStudy (lightn1ng):

where did you get 21 from?

OpenStudy (anonymous):

yea how

OpenStudy (anonymous):

wait is it 9x+3 and y= 21

OpenStudy (0487308):

You can add matching coefficients with every term of a polynomial, as long as they ae proportional.

OpenStudy (0487308):

Basically, 9x = 3(3x), so I multiplied every other term by three to make the two equations coincident.

OpenStudy (0487308):

Do you get it now?

OpenStudy (anonymous):

OO thanks

OpenStudy (0487308):

No problem, and by the way, it is 9x + 3y = 21, not 9x+3 and y = 21

OpenStudy (0487308):

Any other problems you have you can tag me in ;)